Problem 6.4 Mikey W. Smitty, an emerging rapper, is getting ready to cut his first CD, called “Western Rap.” He is looking for a production company to produce his "Western Rap" CD. Mikey W. Smitty is confident that demand for his “Western Rap” CD will substantially exceed the break even point of 500CDs or $7,500 in sales. Studio A charges $25,500 to record the CD and $27.50 per CD to duplicate it. Studio B charges only $18,000 to record the CD, but $32.50 per CD for duplication. The CD will sell for $57.50. What is the breakeven point for Studio A? (Round answer to 2 decimal places, the tolerance is +/- 1). The breakeven point for Studio A is CDs. Link to Text Compare Studio A and B. At what volume of demand should Studio A be chosen? (Round answer to the nearest whole number, the tolerance is +/- 1).
Answer #1

Studio A:

Fixed costs =Costs to record the CD=25500

Variable cost per CD=cost for duplication=27.50

Selling price per CD=57.50

Break even point =Fixed costs/(Price-variable cost)

=25500/(57.50-27.50)

=25500/30

=850 CDs

Studio B:

Fixed costs =Costs to record the CD=18000

Variable cost per CD=cost for duplication=32.50

Selling price per CD=57.50

Break even point =Fixed costs/(Price-variable cost)

=18000/(57.50-32.50)

=18000/25

=720 CDs

Studio A should be chosen when the value of demand exceeds 850 CDs.(i.e demand for CDs >=850)